The absorbance (?)(A) of a solution is defined as

?=log10(?0?)A=log10⁡(I0I)

where ?0I0 is the incident‑light intensity and ?I is the transmitted‑light intensity. Absorbance is also defined as

?=???A=ϵcl

where ?ϵ is the molar absorption coefficient (extinction coefficient) in units of M−1cm−1,M−1cm−1, ?c is the molar concentration, and ?l is path length in centimeters.

Daniella prepares a 1 mg/ml1 mg/ml myoglobin solution. The molecular weight of myoglobin is 17.8 kDa.17.8 kDa. Given that the ?ϵ of myoglobin is 15,000 M−1cm−1,15,000 M−1cm−1, calculate the absorbance of the myoglobin solution across a 1 cm1 cm path. Calculate your answer to two decimal places.

?=A=

What percentage of the incident light is transmitted by this solution? Calculate your answer to one decimal place.

transmitted percentage =

Can someone help me solve for this and explain to me how to solve it please?
Answer #1

from Lambert - Beer law

A = c l

A is  absorbance

molar extinction coefficient

given , = 15000 M-1 cm-1

l = path length = 1 cm

concentration of myoglobin  = 1 mg/ml

given molecular weight of myoglobin = 17.8 KD = 17800 g/mol

molar concentration (c) =

or, c = = 5.62 *10-5 M

now, Absorbance (A) = 15000 * 5.62 *10-5 * 1 = 0.84 ( 2 decimal point)

transmittance ; T = , and transmittance percentage (% T)  = * 100

now, relation between percentage transmittance and absorbance is

so, A = 2 - log10 (%T)

putting the calculated value of A

0.843 = 2 - log10(%T)

or, log10(%T) = (2 - 0.84) = 1.16

or, %T = 10 1.16 = 14.4 (1 decimal point)

hence 14.4 % incident light is transmitted by the solution